2. Standard license plates in Virginia consist of three letters followed by four numbers. A sample license plate is pictured below: OCT . VIRGINIA . 94 ZTT-7019 Note that the four digit number never begins with zero. llow many license plates are available (a) allowing repctitions and that must contain exactly one C, no Ns, and at least one 4? (b) with no repetitions and without vowels and whose 4 digit number is even?
Answer #1

a) total number wit 1 C; and no N's =N(choose 1 out of 3 place for a C and then there are 24 choices for rest of 2 places for alphabets and for first digit 9 choice and for rest 3 10 choices are there)=3*24*24*9*10*10*10=15552000

total number wit 1 C; and no N's and no 4 =3*24*24*8*9*9*9=10077696

hence number with 1 C; and no N's and at least one 4 =15552000-10077696=5474304

b) for there are 21 alphabets without vowel ; and 5 even numbers (0,2,4,6,8)

hence number with no repetiotions and without vowels and whose 4 digit is even =21*20*19*4*4*3*2=766080